Session: May 1997 (session 2)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti / D. Dallacasa at Tue Jun 10 14:51:39 METDST 1997) Success
3C380 set at 11:10:09 UT. High humidity produced system temperatures slightly higher than standard values.

Torun (K.Borkowski at Tue Jun 10 14:56:06 METDST 1997) Minor failures
No problems.

Torun (K.Borkowski at Tue Jun 10 14:57:55 METDST 1997) Success
Oops! Really no problems.

Noto (C.Trigilio at Wed Jun 11 12:38:58 METDST 1997) Success
No problem reported. Last scan in a new tape for only 3 min (!) at very low elevation. Only tape 1 and 2 sent to Socorro (tape 3 pratically with no data).

Jodrell (Alastair Gunn at Mon Jun 16 11:31:00 METDST 1997) Minor failures
Telescope not on 3C380 until 00:32 UT. No other known problems.

Effelsberg (dgraham at Mon Jun 16 14:34:24 METDST 1997) Success

Onsala (John Conway at Wed Jun 18 13:46:05 METDST 1997) Success
No nkown problem. This expt like all 6cm expts used the 25m antenna (onsala85).

Medicina (Fredrik T Rantakyro at Mon Jul 21 11:20:40 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97 directory with the name GP15B.ANTAB. The file is ready to be used with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific comments regarding this experiment.